Circular No. 7508

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000dt IN UGC 3411 G. M. Hurst, Basingstoke, England, reports the discovery by M. Arm-

strong, Rolvenden, of an apparent supernova (mag 17.5) on unfiltered CCD images (limiting mag 19.5) taken with a 0.35-m Schmidt-Cassegrain telescope in the course of the U.K. Nova/Supernova Patrol on Oct. 13.058, 13.914, and 14.019 UT. SN 2000dt is located at α = 6^{h}10^{m}54^{s}.68, δ = +62^{o}01′27′′.2 (equinox 2000.0), which is 2′′ west and 16′′ north of the center of UGC 3411. The new object appears neither on a CCD image taken by Armstrong on Sept. 21 (limiting mag 19.5), nor on the following Palomar Sky Survey images: 1990 Jan. 1 (red, limiting mag 20.8); 1993 Oct. 12 and Dec. 6 (blue, limiting mag 22.5); and 1996 Oct. 16 and Dec. 2 (red, limiting mag 20.8).

Y. Sato and W. D. Li, University of California at Berkeley, on behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report their independent discovery of SN 2000dt with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on unfiltered images taken on Oct. 14.4 (mag ∼ 17.8) and 15.4 UT (mag ∼ 17.9), providing position end figures 54^{s}.68, 27′′.4 (or 1′′.3 west and 15′′.0 north of the nucleus of UGC 3411). A KAIT image of the same field taken on Mar. 21.1 showed nothing at the position of SN 2000dt (limiting mag ∼ 19.0).

COMETS C/1999 U6, C/2000 T3, C/2000 T4 (SOHO) Further to IAUC 7445 and 7480, D. Hammer reports measurements

for three more Kreutz sungrazing comets. All were visible in the LASCO C2 coronagraph, and C/2000 T4 was also seen in C3 images. C/1999 U6, C/2000 T3, and C/2000 T4 were found by M. Oates, D. Biesecker, and J. Danaher, respectively. The reduced positions and orbits by B. G. Marsden are given on MPEC 2000-T46.

Comet  1999 UT  α_{2000}  δ_{2000}
C/1999 U6  Oct. 17.910  13^{h}20.^{m}8  −10^{o}09′
Comet  2000 UT  α_{2000}  δ_{2000}
C/2000 T3  Oct. 11.254  12^{h}58.^{m}5  − 7^{o}45′
C/2000 T4  12.071  12 58.4  − 8 15
2000 October 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green